# iSEARLE: Improving Textual Inversion for Zero-Shot Composed Image Retrieval

### v1

@@ Abstract
Given a query consisting of a reference image and a relative caption, Composed Image Retrieval (CIR) aims to retrieve target images visually similar to the reference one while incorporating the changes specified in the relative caption. The reliance of supervised methods on labor-intensive manually labeled datasets hinders their broad applicability. In this work, we introduce a new task, Zero-Shot CIR (ZS-CIR), that addresses CIR without the need for a labeled training dataset. We propose an approach named iSEARLE (improved zero-Shot composEd imAge Retrieval with textuaL invErsion) that involves mapping the visual information of the reference image into a pseudo-word token in CLIP token embedding space and combining it with the relative caption. To foster research on ZS-CIR, we present an open-domain benchmarking dataset named CIRCO (Composed Image Retrieval on Common Objects in context), the first CIR dataset where each query is labeled with multiple ground truths and a semantic categorization. The experimental results illustrate that iSEARLE obtains state-of-the-art performance on three different CIR datasets – FashionIQ, CIRR, and the proposed CIRCO – and two additional evaluation settings, namely domain conversion and object composition. The dataset, the code, and the model are publicly available at https://github.com/miccunifi/SEARLE.
@@ I Introduction
When provided with a query consisting of a reference image and a relative caption, Composed Image Retrieval (CIR) [1, 2] seeks to retrieve target images that visually resemble the reference one while including the modifications described in the relative caption. The bi-modal structure of the query offers users a more precise means of specifying the desired image characteristics. It leverages the strengths of both language-based descriptions and visual features, as certain attributes are more effectively communicated through text, while others are better expressed visually. We provide some query examples in Fig. 3.
@@
Datasets for CIR comprise triplets $(I_{r},T_{r},I_{t})$, each including a reference image, a relative caption, and a target image, respectively. The creation of datasets for CIR is costly, primarily because such data is not readily accessible online and automated generation remains a significant challenge. Consequently, researchers are compelled to undertake labor-intensive manual labeling efforts. The manual process entails the identification of reference and target image pairs and the composition of descriptive captions that outline the differences between them. Carrying out this task is both time-consuming and resource-intensive, particularly when building extensive training datasets.
@@
Current works addressing CIR [3, 4, 5, 6, 7, 2, 8] depend on supervised learning to devise methods able to combine the reference image and the relative caption effectively. For example, [5] proposes a fully supervised two-stage approach based on fine-tuning CLIP encoders and training a combiner network. Despite their promising results, the dependence of current approaches for CIR on expensive manually annotated datasets constrains their scalability and applicability in domains outside those of the training datasets.
@@
Building on the conference version of this work [9], we remove the need for costly labeled training data by introducing a new task: Zero-Shot Composed Image Retrieval (ZS-CIR). In ZS-CIR, the goal is to devise an approach that is capable of merging the information of the reference image and the relative caption without requiring supervised learning.
@@
To address the challenges of ZS-CIR, we propose an approach named iSEARLE  [^1] John Searle is an American philosopher renowned for his work on the philosophy of language and how words denote specific objects. (improved zero-Shot composEd imAge Retrieval with textuaL invErsion) based on the frozen pre-trained CLIP [10] vision-language model. Our method simplifies CIR to standard text-to-image retrieval by mapping the reference image into a learned pseudo-word, which is subsequently appended to the relative caption. The pseudo-word corresponds to a pseudo-word token residing in the CLIP token embedding space. We refer to this mapping process with textual inversion, following the terminology introduced in [11]. iSEARLE involves the pre-training of a textual inversion network – denoted as $\phi$ – on an unlabeled image-only dataset. The pre-training process consists of two stages: an Optimization-based Textual Inversion (OTI) with a GPT-powered regularization loss aimed at generating a set of pseudo-word tokens, and the distillation of their knowledge to $\phi$. Upon completion of the training, the network $\phi$ is capable of carrying out textual inversion in a single forward pass. During inference, when presented with a query $(I_{r},T_{r})$, we employ $\phi$ to predict the pseudo-word corresponding to $I_{r}$ and then concatenate it to $T_{r}$. Afterward, we exploit the CLIP common embedding space to perform text-to-image retrieval. Figure 1 shows the workflow of the proposed approach.
[…]
@@ II Related Work
Composed Image Retrieval CIR is a branch of compositional learning, an area that has been widely explored in various vision and language (V&L) tasks. These include visual question answering [20, 21], image captioning [22, 23, 24], and image synthesis [25, 26, 27]. Compositional learning aims to create joint embedding features that effectively integrate and express information from both the textual and visual domains.
The research on CIR spans several domains, including fashion [12, 13, 14, 15], natural images [16, 2], and synthetic images [1]. The task was first introduced in [1], where the authors propose a residual gating method for composing image-text features, aiming to merge multimodal information effectively. [28] presents a training strategy that merges graph convolutional networks with existing composition methods. [8] consider image style and content separately by using two different neural network modules. More recently, the use of the CLIP model as a backbone for CIR has received increasing attention [4, 5, 3, 6]. [4] shows the effectiveness of combining out-of-the-box CLIP features with a Combiner network. Building on this, [5] introduces a task-specific fine-tuning step for CLIP encoders. Unlike the aforementioned approaches, the proposed method does not require supervision and uses unlabeled images for training, effectively learning to combine multimodal information without relying on a manually annotated CIR dataset.
@@
Textual Inversion In text-to-image synthesis, the idea of mapping images to a single pseudo-word is emerging as a powerful technique for generating personalized images [11, 39, 40, 41]. [11] performs textual inversion by relying on the reconstruction loss of a latent diffusion model [25]. Additionally, [39] also fine-tunes a pre-trained text-to-image diffusion model.
Besides personalized text-to-image synthesis, textual inversion has also been applied to image retrieval tasks [42, 18, 43, 37, 36]. Specifically, PALAVRA [42] addresses personalized image retrieval by pre-training a mapping function and then optimizing the predicted pseudo-word token at inference time. Several works employ textual inversion to address ZS-CIR [18, 37, 36]. LinCIR [37] is a language-only approach for training the textual inversion network. Context-I2W [36] is based on a transformer-based textual inversion network trained on the image-caption pairs of the CC3M dataset [44]. Moreover, the textual inversion process is dependent also on the query text. However, this comes at the cost of requiring a double forward pass of the text encoder and a more complex network architecture than the proposed single MLP approach. The method most similar to ours is Pic2Word [18]. Pic2Word relies on a textual inversion network trained on the 3M images of CC3M using only a cycle contrastive loss. In contrast, we train our textual inversion network on only 3% of the data and use a weighted sum of distillation and regularization losses. The distillation loss leverages the information provided by a set of pre-generated tokens obtained via optimization-based textual inversion.
Knowledge Distillation Knowledge distillation is a machine learning technique in which a simple model (referred to as the student) learns to replicate the behavior of a more complex one (referred to as the teacher) by learning from its predictions [45]. This method has proven effective in various computer vision tasks, such as image classification [45, 46, 47], object detection [48, 49] and text-to-image synthesis [27, 50], improving model compression, computational efficiency, and accuracy. In our work, we refer to knowledge distillation as the process of transferring knowledge from a computationally expensive optimization method (teacher) to a more efficient neural network (student). Specifically, we train a textual inversion network to emulate the output of an optimization-based textual inversion using a distillation loss. From a different point of view, our lightweight network can be viewed as a surrogate model of the more resource-intensive optimization technique.
@@ III Proposed Approach
Figure 2: Overview of our approach. Left: we generate a pseudo-word token $v_{*}$ from an image $I$ with an iterative Optimization-based Textual Inversion (OTI). We force $v_{*}$ to represent the image content with a cosine loss $\mathcal{L}_{content}$. We add Gaussian noise to the text features before computing $\mathcal{L}_{content}$. We assign a concept word to $I$ with a CLIP zero-shot classification and feed the prompt “a photo of {concept}” to GPT to continue the phrase, resulting in $\widehat{T}$. Let $S_{*}$ be the pseudo-word associated with $v_{*}$. We craft $\widehat{T}_{*}$ by replacing in $\widehat{T}$ the concept with $S_{*}$. $\widehat{T}$ and $\widehat{T}_{*}$ are then employed for a contextualized regularization with $\mathcal{L}_{gpt}$. Right: we train a textual inversion network $\phi$ on unlabeled images. Given a set of pseudo-word tokens pre-generated with OTI, we distill their knowledge to $\phi$ through a contrastive loss $\mathcal{L}_{distil}$. We regularize the output of $\phi$ with the same GPT-powered loss $\mathcal{L}_{gpt}$ employed in OTI and an additional penalty term $\mathcal{L}_{pen}$. B represents the number of images in a batch.
@@
Our objective is dual. Firstly, the pseudo-word token has to accurately capture the content of the reference image. In other words, the text features related to a basic prompt comprising the pseudo-word should closely align with the corresponding image features. Secondly, such a pseudo-word should be able to effectively integrate and communicate with the text of the relative caption. While theoretically a single image could be mapped to multiple pseudo-word tokens, in this work, we opt for using a single one. This decision is based on findings from [11], which shows that a single pseudo-word token is sufficient to encode the information of an image effectively.
@@
iSEARLE entails pre-training a textual inversion network $\phi$ using an unlabeled image-only dataset. This training process unfolds in two distinct stages. Firstly, we rely on an Optimization-based Textual Inversion (OTI) method, which iteratively produces a set of pseudo-word tokens by exploiting a GPT-based regularization loss. Secondly, we train $\phi$ via knowledge distillation from the pre-generated pseudo-word tokens. The $\phi$ network outputs the pseudo-word token associated with an image in a single forward pass by taking as input its visual features, previously extracted via the CLIP image encoder.
@@
At inference time, the input for CIR is given by a query $(I_{r},T_{r})$ corresponding respectively to the reference image and the relative caption. We generate the pseudo-word token $v_{*}$ associated with the reference image as $v_{*}=\phi(I_{r})$. Let $S_{*}$ be the pseudo-word related to $v_{*}$. To effectively integrate the visual information of $I_{r}$ with $T_{r}$, we build the template “a photo of $S_{*}$ that {relative caption}” and extract its features via the CLIP text encoder. Note that these text features provide a multimodal representation of the reference image and its associated relative caption, as they encompass both textual and visual information. Finally, we carry out standard text-to-image retrieval within an image database using the extracted text features. We provide an overview of the workflow of our method in Fig. 1.
@@
Fundamentally, both OTI and $\phi$ carry out the same operation, i.e. mapping the visual features of an image into a pseudo-word token by means of a textual inversion. Consequently, one could directly utilize OTI at inference time without the need for $\phi$. However, $\phi$ offers considerably improved efficiency compared to OTI, which needs a non-negligible amount of time to be performed. Considering that OTI has demonstrated its effectiveness in generating fruitful pseudo-word tokens (see Sec. V), we propose to distill their knowledge into a feed-forward network. Our approach strives to maintain the powerful expressiveness of OTI while achieving a negligible inference time. From now on, we refer to our approach as iSEARLE when relying on $\phi$ for generating the pseudo-word token and as iSEARLE-OTI when we directly utilize OTI for inference.
@@ III-A Optimization-based Textual Inversion (OTI)
First, we randomly initialize the pseudo-word token $v_{*}$ and associate the pseudo-word $S_{*}$ with it. Then, we craft a template sentence $T$ such as “a photo of $S_{*}$” and process it with the CLIP text encoder $\psi_{T}$, resulting in $y=\psi_{T}(T)$. Following [42], we randomly sample $T$ from a given set of templates. We employ the CLIP image encoder $\psi_{I}$ to extract the visual features $x=\psi_{I}(I)$.
Our goal is to obtain a pseudo-word token $v_{*}$ that captures the informative content of $I$. To this end, in our preliminary work [9], we directly minimize the discrepancy between the image and text features by leveraging CLIP common embedding space. However, [19] shows that in vision-language models such as CLIP the features associated with text and images correspond to different regions of the joint embedding space. In other words, text and image embeddings fall into separate clusters in the feature space. This phenomenon is commonly referred to as modality gap [19, 51]. To mitigate this issue, [51] proposes a simple and training-free strategy that involves adding Gaussian noise to the text features. Intuitively, the noise reduces the modality gap by spreading out the text embeddings to make them overlap with the image ones.
Inspired by [51], we propose to add Gaussian noise to the text features $y$ before minimizing their discrepancy to the image features $x$. Specifically, we compute $\overline{y}=y+n$, where $n\sim\mathcal{N}(0,\gamma^{2})$ is a sample drawn from a Gaussian distribution with variance $\gamma^{2}$. Finally, we employ a cosine CLIP-based loss to maximize the similarity between the image and noisy text features:
@@
Therefore, differently from our preliminary work [9], we mitigate the modality gap issue before computing the loss, resulting in improved performance (see Sec. V-C). In addition, despite addressing the modality gap problem, [51] does not directly contrast noisy text features with image ones, as the authors train their model without relying on visual data. On the contrary, we show that adding Gaussian noise to the text features is effective also when they are directly compared to the image features in the loss computation.
Relying solely on $\mathcal{L}_{content}$ is inadequate for generating a pseudo-word capable of interacting with other words of the CLIP dictionary. Indeed, similar to [42], we observe that $\mathcal{L}_{content}$ pushes the pseudo-word token into sparse regions of CLIP token embedding space that differ from those encountered during CLIP’s training. This phenomenon, akin to effects observed in GAN inversion works [52, 53], hampers the ability of the pseudo-word token to communicate effectively with other tokens. To address this limitation, we propose a novel regularization technique that constrains the pseudo-word token to reside on the CLIP token embedding manifold, thereby enhancing its reasoning capabilities. Relying on CLIP zero-shot capabilities, we carry out a zero-shot classification of the image $I$. To classify the images we employ a vocabulary originating from the $\sim$20K class names of the Open Images V7 dataset [54]. Specifically, we assign the $k$ most similar distinct class names to each image, with $k$ being a hyperparameter. We will refer to these class names as concepts, so, in other words, we associate each image to $k$ different concepts. Differently from [42], we do not require the concepts as input.

### v2

@@ Abstract
Given a query consisting of a reference image and a relative caption, Composed Image Retrieval (CIR) aims to retrieve target images visually similar to the reference one while incorporating the changes specified in the relative caption. The reliance of supervised methods on labor-intensive manually labeled datasets hinders their broad applicability to CIR. In this work, we introduce a new task, Zero-Shot CIR (ZS-CIR), that addresses CIR without the need for a labeled training dataset. We propose an approach, named iSEARLE (improved zero-Shot composEd imAge Retrieval with textuaL invErsion), that involves mapping the visual information of the reference image into a pseudo-word token in the CLIP token embedding space and combining it with the relative caption. To foster research on ZS-CIR, we present an open-domain benchmarking dataset named CIRCO (Composed Image Retrieval on Common Objects in context), the first CIR dataset where each query is labeled with multiple ground truths and a semantic categorization. The experimental results illustrate that iSEARLE obtains state-of-the-art performance on three different CIR datasets – FashionIQ, CIRR, and the proposed CIRCO – and two additional evaluation settings, namely domain conversion and object composition. The dataset, code, and model are publicly available at https://github.com/miccunifi/SEARLE.
@@ I Introduction
When provided with a query consisting of a reference image and a relative caption, Composed Image Retrieval (CIR) [1, 2] seeks to retrieve target images that visually resemble the reference one while including the modifications described in the relative caption. The bi-modal structure of the query allows users to specify the desired image characteristics more precisely. It leverages the strengths of both language-based descriptions and visual features, as certain attributes are more effectively communicated through text, while others are better expressed visually. We provide some query examples in Fig. 3.
@@
Datasets for CIR comprise triplets $(I_{r},T_{r},I_{t})$, each including a reference image, a relative caption, and a target image, respectively. The creation of datasets for CIR is costly, primarily because such data is not readily accessible online, and automated generation remains a significant challenge. Consequently, researchers are compelled to undertake labor-intensive manual labeling efforts. The manual process entails the identification of reference and target image pairs and the composition of descriptive captions that outline the differences between them. Carrying out this task is both time-consuming and resource-intensive, particularly when building extensive training datasets.
@@
Current works addressing CIR [3, 4, 5, 6, 7, 2, 8] depend on supervised learning to devise methods able to combine the reference image and the relative caption effectively. For example, [5] proposes a fully supervised two-stage approach based on fine-tuning CLIP encoders and training a combiner network. Despite their promising results, the dependence of current CIR approaches on expensive manually annotated datasets constrains their scalability and applicability in domains outside those of the training datasets.
@@
Building on the conference version of this work [9], we remove the need for costly labeled training data by introducing a new task: Zero-Shot Composed Image Retrieval (ZS-CIR). In ZS-CIR, the goal is to devise an approach capable of merging the information of the reference image and the relative caption without requiring supervised learning.
@@
To address the challenges of ZS-CIR, we propose an approach named iSEARLE  [^1] John Searle is an American philosopher renowned for his work on the philosophy of language and how words denote specific objects. (improved zero-Shot composEd imAge Retrieval with textuaL invErsion) based on the frozen pre-trained CLIP [10] vision-language model. Our method simplifies CIR, casting it to a standard text-to-image retrieval task by mapping the reference image into a learned pseudo-word, which is subsequently appended to the relative caption. The pseudo-word corresponds to a pseudo-word token residing in the CLIP token embedding space. We refer to this mapping process with textual inversion, following the terminology introduced in [11]. iSEARLE involves the pre-training of a textual inversion network – denoted as $\phi$ – on an unlabeled image-only dataset. The pre-training process consists of two stages: an Optimization-based Textual Inversion (OTI) with a GPT-powered regularization loss aimed at generating a set of pseudo-word tokens, and the distillation of their knowledge to $\phi$. Upon completion of the training, the network $\phi$ is capable of carrying out textual inversion in a single forward pass. During inference, when presented with a query $(I_{r},T_{r})$, we employ $\phi$ to predict the pseudo-word corresponding to $I_{r}$ and then concatenate it to $T_{r}$. Afterward, we exploit the CLIP common embedding space to perform text-to-image retrieval. Figure 1 shows the workflow of the proposed approach.
The majority of existing CIR datasets focus on specific domains, such as fashion [12, 13, 14, 15], birds [16], or synthetic objects [1]. To the best of our knowledge, the CIRR dataset [2] stands alone in encompassing natural images within an open domain. However, CIRR suffers from two main problems. Firstly, it includes numerous false negatives, potentially leading to imprecise performance evaluations. Secondly, the queries often neglect the visual content of the reference image, rendering the task addressable through standard text-to-image techniques, as shown by the results reported in Tab. III. Additionally, existing CIR datasets provide only a single annotated ground truth image per query. To address these shortcomings and foster research on ZS-CIR, we introduce an open-domain benchmarking dataset named CIRCO [^2] CIRCO is pronounced as /irko/. (Composed Image Retrieval on Common Objects in context). CIRCO comprises validation and test sets derived from images within the COCO dataset [17]. As a benchmarking dataset for ZS-CIR, a large training set is not needed, leading to a considerable reduction in labeling effort. To overcome the single ground truth limitation of existing CIR datasets, we propose to leverage our method to ease the annotation process of multiple ground truths. Consequently, CIRCO is the first CIR dataset with multiple annotated ground truths, enabling a more comprehensive evaluation of CIR models. In addition, contrary to existing CIR datasets, we provide a semantic categorization of the queries that allows a fine-grained semantic analysis of the results. We release only the validation set ground truths of CIRCO and host an evaluation server, enabling researchers to get performance metrics on the test set [^3] Accessible at: https://circo.micc.unifi.it.
@@
The remainder of this paper is organized as follows. Sec. II reviews related work. Sec. III details our proposed approach. Sec. IV describes the proposed CIRCO dataset. Sec. V presents experimental results and analysis. Sec. VI concludes the paper with final remarks.
@@ II Related Work
Composed Image Retrieval CIR is a branch of compositional learning, an area that has been widely explored in various vision and language tasks. These include visual question answering [20, 21], image captioning [22, 23], and image synthesis [24, 25]. Compositional learning aims to create joint embedding features that effectively integrate and express information from both the textual and visual domains.
The research on CIR spans several domains, including fashion [12, 13, 14, 15], natural images [16, 2], and synthetic images [1]. The task was first introduced in [1], where the authors propose a residual gating method for composing image-text features, aiming to merge multimodal information effectively. More recently, the use of the CLIP model as a backbone for CIR has received increasing attention [4, 5, 3, 6]. [4] shows the effectiveness of combining out-of-the-box CLIP features with a Combiner network. Building on this, [5] introduces a task-specific fine-tuning step for CLIP encoders. Unlike the aforementioned approaches, the proposed method does not require supervision and uses unlabeled images for training, effectively learning to combine multimodal information without relying on a manually annotated CIR dataset.
@@
Textual Inversion In text-to-image synthesis, mapping images to a single pseudo-word is emerging as a powerful technique for generating personalized images [11, 36, 37]. [11] performs textual inversion by relying on the reconstruction loss of a latent diffusion model [24]. Additionally, [36] also fine-tunes a pre-trained text-to-image diffusion model.
Besides personalized text-to-image synthesis, textual inversion has also been applied to image retrieval tasks [38, 18, 39, 34, 33]. Specifically, PALAVRA [38] addresses personalized image retrieval by pre-training a mapping function and then optimizing the predicted pseudo-word token at inference time. Several works employ textual inversion to address ZS-CIR [18, 34, 33]. LinCIR [34] is a language-only approach for training the textual inversion network. Context-I2W [33] is based on a transformer-based textual inversion network trained on the image-caption pairs of the CC3M dataset [40]. Moreover, the textual inversion process is also dependent on the query text. This comes at the cost of requiring a double forward pass of the text encoder and a more complex network architecture than the proposed single MLP approach. The method most similar to ours is Pic2Word [18]. Pic2Word relies on a textual inversion network trained on the 3M images of CC3M using only a cycle contrastive loss. In contrast, we train our textual inversion network on only 3% of the data and use a weighted sum of distillation and regularization losses. The distillation loss leverages the information provided by a set of pre-generated tokens obtained via optimization-based textual inversion.
Knowledge Distillation Knowledge distillation is a machine learning technique in which a simpler model (the student) learns to replicate the behavior of a more complex one (the teacher) by learning from its predictions [41]. This method has proven effective in various computer vision tasks, such as image classification [41, 42, 43], object detection [44, 45], and text-to-image synthesis [25, 46], improving model compression, computational efficiency, and accuracy. In our work, we refer to knowledge distillation as the process of transferring knowledge from a computationally expensive optimization method (teacher) to a more efficient neural network (student). Specifically, we train a textual inversion network to emulate the output of an optimization-based textual inversion using a distillation loss. From a different point of view, our lightweight network can be viewed as a surrogate model of the more resource-intensive optimization technique.
@@ III Proposed Approach
Figure 2: Overview of our approach. Left: single iteration of the iterative Optimization-based Textual Inversion (OTI) method to generate a pseudo-word token $v_{*}$ from an image $I$. We force $v_{*}$ to represent the image content with a cosine loss $\mathcal{L}_{content}$. We add Gaussian noise to the text features before computing $\mathcal{L}_{content}$. We assign a concept word to $I$ with a CLIP zero-shot classification and feed the prompt ``a photo of {concept}" to GPT to continue the phrase, resulting in $\widehat{T}$. Let $S_{*}$ be the pseudo-word associated with $v_{*}$. We craft $\widehat{T}_{*}$ by replacing in $\widehat{T}$ the concept with $S_{*}$. $\widehat{T}$ and $\widehat{T}_{*}$ are then employed for a contextualized regularization with $\mathcal{L}_{gpt}$. Right: pre-training of textual inversion network $\phi$ on unlabeled images. Given a set of pseudo-word tokens pre-generated with OTI, we distill their knowledge to $\phi$ through a contrastive loss $\mathcal{L}_{distil}$. We regularize the output of $\phi$ with the same GPT-powered loss $\mathcal{L}_{gpt}$ employed in OTI and an additional penalty term $\mathcal{L}_{pen}$. B represents the number of images in a batch.
@@
Our objective is dual. First, the pseudo-word token must accurately capture the content of the reference image. In other words, the text features related to a basic prompt comprising the pseudo-word should closely align with the corresponding image features. Second, the pseudo-word must effectively integrate and communicate with the text of the relative caption. While a single image can be mapped to multiple pseudo-word tokens, we opt for using a single one, as it proves to be sufficient to encode the information of an image effectively (see Tab. XI). Moreover, from our preliminary experiments, using a single pseudo-word token achieves better performance than relying on multiple ones, as also observed by [11].
@@
iSEARLE entails pre-training a textual inversion network $\phi$ using an unlabeled image-only dataset through a two-stage process. First, we rely on an Optimization-based Textual Inversion (OTI) method, which iteratively produces a set of pseudo-word tokens by exploiting a GPT-based regularization loss. Second, we train $\phi$ via knowledge distillation from the pre-generated pseudo-word tokens. The $\phi$ network outputs the pseudo-word token associated with an image in a single forward pass by taking as input its visual features, previously extracted via the CLIP image encoder.
@@
At inference time, the input for CIR is given by a query $(I_{r},T_{r})$ corresponding respectively to the reference image and the relative caption. We generate the pseudo-word token $v_{*}$ associated with the reference image as $v_{*}=\phi(I_{r})$. We denote by $S_{*}$ the pseudo-word corresponding to the pseudo-word token $v_{*}$, i.e. the counterpart of $v_{*}$ expressed in natural language. To effectively integrate the visual information of $I_{r}$ with $T_{r}$, we build the template ``a photo of $S_{*}$ that {relative caption}" and extract its features via the CLIP text encoder. Note that these text features provide a multimodal representation of the reference image and its associated relative caption, as they encompass both textual and visual information. Finally, we carry out standard text-to-image retrieval within an image database using the extracted text features. We provide an overview of the workflow of our method in Fig. 1.
@@
Fundamentally, both OTI and $\phi$ carry out the same operation, i.e. mapping the visual features of an image into a pseudo-word token by means of a textual inversion. Consequently, one could directly utilize OTI at inference time without the need for $\phi$. However, $\phi$ offers considerably improved efficiency compared to OTI, which needs a non-negligible amount of time to be performed (see Sec. V-A). Considering that OTI has demonstrated its effectiveness in generating fruitful pseudo-word tokens (see Sec. V), we propose to distill their knowledge into a feed-forward network. Our approach strives to maintain the powerful expressiveness of OTI while achieving a negligible inference time. From now on, we refer to our approach as iSEARLE when relying on $\phi$ for generating the pseudo-word token and as iSEARLE-OTI when we directly utilize OTI for inference.
@@ III-A Optimization-based Textual Inversion (OTI)
First, we randomly initialize the pseudo-word token $v_{*}$ and associate the pseudo-word $S_{*}$ with it. Then, we craft a template sentence $T$, such as ``a photo of $S_{*}$", and process it with the CLIP text encoder $\psi_{T}$, resulting in $y=\psi_{T}(T)$. Following [38], we randomly sample $T$ from a given set of templates. We employ the CLIP image encoder $\psi_{I}$ to extract the visual features $x=\psi_{I}(I)$.
Our goal is to obtain a pseudo-word token $v_{*}$ that captures the informative content of $I$. To this end, in our preliminary work [9], we directly minimize the discrepancy between the image and text features by leveraging the CLIP common embedding space. However, [19] shows that in vision-language models such as CLIP the features associated with text and images correspond to different regions of the joint embedding space. In other words, text and image embeddings fall into separate clusters in the feature space. This phenomenon is commonly referred to as modality gap [19, 47, 48]. To mitigate this issue, [48] proposes a simple and training-free strategy that involves adding Gaussian noise to the text features. Intuitively, the noise reduces the modality gap by spreading out the text embeddings to make them overlap with the image ones.
Inspired by [48], we propose to add Gaussian noise to the text features $y$ before minimizing their discrepancy with the image features $x$. Specifically, we compute $\overline{y}=y+n$, where $n\sim\mathcal{N}(0,\gamma^{2})$ is drawn from a Gaussian distribution with variance $\gamma^{2}$. Finally, we employ a cosine loss to maximize the similarity between the image and noisy text features:
@@
Therefore, differently from our preliminary work [9], we mitigate the modality gap issue before computing the loss, resulting in improved performance (see Sec. V-C). In addition, despite addressing the modality gap problem, [48] does not directly contrast noisy text features with image ones, as the authors train their model without relying on visual data. On the contrary, we show that adding Gaussian noise to the text features is effective even when they are directly compared to the image features in the loss computation.
Relying solely on $\mathcal{L}_{content}$ is inadequate for generating a pseudo-word capable of interacting with other words of the CLIP dictionary. Indeed, similar to [38], we observe that $\mathcal{L}_{content}$ pushes the pseudo-word token into sparse regions of CLIP token embedding space that differ from those encountered during CLIP's training. This phenomenon, akin to effects observed in GAN inversion works [49, 50], hampers the ability of the pseudo-word token to communicate effectively with other tokens. To address this limitation, we propose a novel regularization technique that constrains the pseudo-word token to reside on the CLIP token embedding manifold, thereby enhancing its interaction capabilities. Relying on CLIP zero-shot capabilities, we carry out a zero-shot classification of the image $I$. To classify the images, we employ a vocabulary originating from the $\sim$20K class names of the Open Images V7 dataset [51]. Specifically, we assign the $k$ most similar distinct class names to each image, with $k$ being a hyperparameter. We will refer to these class names as concepts, so, in other words, we associate each image to $k$ different concepts. Differently from [38], we do not require the concepts as input.